5.6L of gas at STP weighs equal to 8g.The vapor density of the gas is
1
Expert's answer
2012-07-10T07:42:54-0400
STP = "standard temperature and pressure" as a temperature of 273.15 K (0 °C, 32
°F) and an absolute pressure of 100 kPa (14.504 psi, 0.986 atm, 1 bar). R =
8.314 L*kPa / K*mol
pV = nRT = mRT/M
Vapour density is the density of a
vapour in relation to that of hydrogen. Therefore:
vapour density = molar
mass of gas / molar mass of H2 = molar mass of gas / 2.0016
Molar mass is

M = mRT / pV = 8g * 8.314L*kPa/K*mol * 273.15K / 100kPa * 5.6L = 32
g/mol
vapour density = 32g/mol / 2g/mol = 16
The vapor density of the gas
is 16.